# The precision measurement of the electron anti-neutrino spectrum in beta-decay of nuclei

### v1

@@ Abstract
The 144Ce – 144Pr electron antineutrino source is one of the most suitable radiochemical sources for experiments searching for the oscillations of active neutrinos to the light sterile state. In the current work the -spectra of – source have been measured by two types of -spectrometers based on silicon Si(Li) detectors in order to determine the energy spectrum of electron anti-neutrino emitted in the -decay of nuclei. The nuclear form factor of the ground state beta transition in has been obtained with high precision: . The reduced cross section for the inverse beta decay reaction on hydrogen for the electron anti-neutrino source has been defined as that provides sufficient sensitivity to search for a sterile neutrino with a mass of eV and a mixing angle using standard disappearance method.
@@ I INTRODUCTION
The “reactor antyneutrino anomaly” (RAA) has shown a discrepancy between the expected (based on a new anti-neutrino flux calculations [11, 12]) and the observed count rates of inverse beta-decay (IBD) reaction at the detectors installed close, m, to the nuclear reactors. This effect has been explained via mixing between electron and sterile (anti)neutrino with oscillation parameters and [13, 14]. However, the recent results of reactor experiments [15, 16, 17], as well as the new data on the spectrum of reactor neutrinos [18] probably do not confirm the existence of the RAA and its oscillation solution [19].
@@
The expected sensitivity of the experiment with source relative to the sterile neutrino oscillation parameters depends on the accuracy of flux determination as well as the energy spectrum shape for the electron anti-neutrino emitted in the decay of nuclei. The purpose of this work was a precise measurement of the spectra from source and determination of the spectrum of electron anti-neutrinos emitted in the decay nuclei with great precision.
@@ II EXPERIMENTAL SETUP
The spectra of the source were measured by two setups: one using a beta-spectrometer in the classical ”target - detector” configuration, and the other relies an originally developed total absorption -spectrometer consisting of a pair of Si(Li)-detectors. These measurements correspond to the application of two approaches to determining the detector response function - accurate M-C modeling and the use of a detector response function close to a Gaussian function.
@@ II.1 Spectrometer “target-detector” type
The Si(Li)-detector produced from p-type single-crystal silicon with a resistivity of k and a carrier lifetime of s using standard technology that has been developed and tested by the Petersburg Nuclear Physics Institute. The characteristics of such detectors were described in [33, 34, 35].
@@
The 144Ce source in the form of a dried drop of colloidal solution was placed on the surface of a maylar substrate with thickness of m. The substrate was fixed in a delrin ring and placed on a supporting structure at the distance of mm from the detector surface. The thickness distribution over the spot of the dried source 144Ce was determined from the measured energy spectrum of alpha particles originating from the decays of 241Am and 244Cm, which were preserved in the source as low activity impurities after purification of the 144Ce source. The average thickness of the source was determined as m.
The entire installation of the Si(Li)-detector and source was placed in a vacuum cryostat and was cooled down to the temperature of liquid nitrogen (LN).
The Si(Li)-detector was equipped with a charge-sensitive preamplifier with a field-effect transistor placed inside the vacuum cryostat and cooled down to the LN temperature. The negative bias voltage of kV was applied directly to the gold coating of the detector.
@@
The energy calibration was performed with help of total absorption peaks for gamma quanta with energies of keV and keV of the 207Bi source (Fig. 2). The energy resolution FWHM (full width at half maximum), determined by the conversion electron peak with an energy of keV, was keV ( eV).
Measurements were carried out continuously for 2024 hours, with short technological stops. The long duration of the measurements was defined by the low activity of the 144Ce source, which was Bk. The drift of the gain coefficient of the Si(Li) detector spectrometric channel during the long-term measurements was compensated by re-calibrations driven by the peak of conversion electrons with an energy of 91.5 keV at decays of 144Ce and the 5.902 MeV -peak of 244Cm, which were presented in the measured spectra.
@@
The original spectrometer with 4 geometry was designed on the basis of two Si(Li) detectors with the thicknesses of about mm each, which exceeds the ionisation path of electrons with an energy of MeV. The detectors were specially manufactured for this experiment and they were made in the “top-hat” shape with external diameters of mm and mm, height of mm and diameters of the sensitive area of mm and mm. The Si(Li) detectors had different external diameters for the convenience of assembling the spectrometer with two detectors docked closely (Fig. 1, (B)) [36, 37, 38].
@@
The total registered energy spectrum is obtained as the sum of the single event spectra from each of the Si(Li)-detectors and the total energy spectrum of events, recorded by both detectors in coincidence. The obtained spectrum solves the response function problem associated with backscattering of electrons from the detector surface.
@@ III DATA ANALYSIS
Decays of the to the ground ( keV, %) and first excited ( keV, %) states of the nucleus correspond to first-order forbidden non-unique and unique transitions, respectively. Decay into the level of with an energy of keV ( keV, %) is an allowed transition, the fitting of the beta spectrum does not require additional shape corrections as in the case of forbidden transitions [39].
The main objective of the work is to determine the spectrum of electron antineutrinos in the decay of with an energy above keV, the threshold of the inverse beta decay reaction on hydrogen. This is possible at electron energies less than keV in the case of a transition to the ground state and at an electron energy less than keV in the decay to the first excited state.
@@
The peculiarity of the described above -spectrometer is that it registers with high efficiency the IB photons with energy up to 100 keV, which are considered as part of the radiative corrections. For example, 1 MeV electron appearing in -decay emits 75% of the IB photons with an energy of less than 100 keV, of which only 6% leave the -detector without interaction. As a result, we used two different forms of the radiative correction function fitting the data of ”target-detector” and -spectrometer. For the -spectrometer we used the correction introduced for the neutrino spectrum [48, 49] with the natural replacement of by .
@@ III.1 Analysis of the data of spectrometer “target-detector” type
The nuclear shape-factor is the function searched in the problem of the -spectrum measurement. The form factor was parameterized in powers of similarly to [51]:
where and are free parameters. As a result of the analysis, the following expression for the form factor function was obtained:
The accuracy of the obtained result (3) is limited by the weak activity of the 144Ce source, which requires a small distance from the detector to the target, that leads to large angles of incidence of electrons onto the crystal surface and an increased proportion of backscattered electrons. The accuracy in determination of the the nuclear form factor parameters and could be significantly increased by usage of a more active source with “target-detector” spectrometer.
@@ III.2 Analysis of the data of 4-spectrometer
Although the response of the spectrometer is close to the Gaussian function, one should take into account the passage of the insensitive layers of the source and detectors. In order to account for these effects detailed M-C calculations were carried out for different electron energies.
The detector response function determined as a result of fitting looks like as a convolution of a Gaussian function with an exponential tail, which describes small energy losses in the target and non-sensitive layers of the detectors. As an example, for 1 MeV electrons the standard deviation for response function (the square root of the second central moment ) is 56.8 keV and the fraction of electrons detected with an energy less then 0.850 MeV is 3.4%.
The response function calculated by the M-C method was described analytically as:
@@
As a result of the analysis, the following values of the parameters and of the nuclear form factor function were obtained:
@@
The -spectrum has been previously measured in several works [51, 52, 53, 54, 55] that used different expressions for the Fermi function and for the form factor. Therefore, a direct comparison of the results obtained is difficult. Nevertheless, our result for is more precise but consistent with [55, 54] within experimental uncertainties.
@@ IV NEUTRINO SPECTRUM
Fig. 7 illustrates the spectrum of electron antineutrinos from 144Pr source. The numerical values for neutrinos with energies above the IBD threshold are shown in the Tab. 1. The uncertainties of are calculated with toy Monte-Carlo method based on the uncertainties of the evaluated form-factor parameters and from (5) and taking into account their correlation. The parameters evaluated with this method were simulated 10000 times and used for performing 2-dimensional spectrum density plot. The RMS of density distributions of its monoenergetic projections was taken as the statistical uncertainty of the spectrum.
@@
Neutrinos from the decay of have the same intensity as -neutrinos, but their energies are less than 318 keV (the region to the left of line 1 in Fig. 7). The threshold value of the IBD reaction on hydrogen in the laboratory system is 1.806 MeV (line 2).
@@
Considering the precision given by our measurement one could apply toy Monte-Carlo to convert the uncertainties on the form-factor parameters to the fraction of the -spectrum above the threshold IBD reaction that is so the statistic precision of this value is as low as 0.01 %. The systematic uncertainty evaluated as the difference of the values derived for two boundary cases of low-energy screening correction is 0.00022 (0.044 %). Greater systematic error is associated with the uncertainty of the energy scale at 1.806 MeV that lead to 0.0065 (0.13 %).
@@
The expected number of IBD events is determined by the integral of the product of the reaction cross section and the neutrino flux (Fig.8). Using the cross section of IBD reaction [56, 57, 58, 59], one could evaluate the experimental neutrino count rate, that is cm2 per one decay of nucleus. The errors reported here are due only to uncertainties in the neutrino spectrum of connected with the shape (mainly due to the uncertainty in the screening correction 0.0041 (0.09%) and energy scale . The latter can be substantially improved by any refinement of the -value of -decay. The obtained precision of in % would correspond to an experimental uncertainty of neutrino count rate expected for a sterile neutrino experiment using Pr source. In this case this value would be sub-dominated by other experimental uncertainties, e.g. the precision of the source calorimetry for Borexino-SOX experiment has reached precision of % [30].
@@
We can also provide an analytical interpolation of the spectrum, that could be expressed in two energy intervals. For neutrino energies above the unique forbidden transition endpoint of MeV, the spectrum is interpolated as follows:
@@ V CONCLUSION
The -spectra of the source were studied using two spectrometers based on semiconductor Si(Li) detectors. The spectra were measured using a beta spectrometer in the classical “target-detector” scheme and with an original full absorption spectrometer consisting of two of Si(Li) detectors. The models of the beta spectrometer response function calculated through the Monte Carlo simulation were tested by fitting the spectrum of the allowed beta transition and have shown good agreement between the response models and experimental data. The function of the nuclear form factor of the ground state beta transition in has been defined as . The electron antineutrino spectrum of the decays was obtained from the measured -spectrum and the form factor parameters. The reduced cross section for the inverse beta decay reaction for the electron anti-neutrino source is cm2 per which is the most accurate measurement up to date.

### v2

# Electron-antineutrino spectrum from precision measurements of -decay

@@ Abstract
The 144Ce – 144Pr electron antineutrino source is one of the most suitable radiochemical sources for experiments searching for the oscillations of active neutrinos to the light sterile state. In the current work the -spectra of – source have been measured by two types of -spectrometers based on silicon Si(Li) detectors in order to determine the energy spectrum of electron anti-neutrino emitted in the -decay of nuclei. The nuclear shape factor of the ground state beta transition in has been obtained with high precision: . The reduced cross section for the inverse beta decay reaction on hydrogen for the electron anti-neutrino source has been defined as that provides sufficient sensitivity to search for a sterile neutrino with a mass of eV and a mixing angle using standard disappearance method.
@@ I INTRODUCTION
The “reactor antineutrino anomaly” (RAA) [11] has shown a discrepancy between the expected (based on a new anti-neutrino flux calculations [12, 13]) and the observed count rates of inverse beta-decay (IBD) reaction at the detectors installed close, m, to the nuclear reactors. This effect has been explained via mixing between electron and sterile (anti)neutrino with oscillation parameters and [14, 15].
However, the recent reactor experiments [16, 17, 18] and new models of reactor neutrino spectra [19], based in particular on the so-called summation approach, made the discrepancy between measurements and theoretical predictions less significant [20, 21].
@@
Decays of the to the ground ( keV, %) and first excited ( keV, %) states of the nucleus correspond to first-order forbidden non-unique and unique transitions, respectively. Decay into the level of with an energy of keV ( keV, %) is an allowed transition, the fitting of the beta spectrum does not require additional shape corrections as in the case of forbidden transitions.
The main objective of -decay study is to determine the spectrum of electron antineutrinos in the decay of with an energy above keV, the threshold of the inverse beta decay reaction on hydrogen. This is possible at electron energies less than keV in the case of a transition to the ground state and at an electron energy less than keV in the decay to the first excited state.
The expected sensitivity of the experiment with source relative to the sterile neutrino oscillation parameters depends on the accuracy of flux determination as well as the energy spectrum shape for the electron anti-neutrino emitted in the decay of nuclei. The sensitivity is evaluated by comparing the observed event rate and positron spectrum of IBD reaction, binned as a function of both energy and distance, with respect to the expected distribution in the presence of oscillations. Three data analyses can be performed: these are the so-called “count rate”, “shape” and “rate + shape” tests. To illustrate the importance of the spectral shape of the antineutrino spectra measurement, just note that the existing shape spectrum uncertainties translates into a % uncertainty on the predicted IBD rate [35, 36, 37, 38].
The purpose of this work was a precise measurement of the spectra from source and determination of the spectrum of electron anti-neutrinos emitted in the decay nuclei with great precision.
@@ II EXPERIMENTAL SETUP
The spectra of the source were measured by two setups: one using a beta-spectrometer in the classical ”target - detector” configuration, and the other relies an originally developed total absorption -spectrometer consisting of a pair of Si(Li)-detectors. These measurements correspond to the application of two approaches to determining the detector response function - accurate Monte Carlo modeling and the use of a detector response function close to a Gaussian function.
@@ II.1 Spectrometer “target-detector” type
The Si(Li)-detector produced from p-type single-crystal silicon with a resistivity of k and a carrier lifetime of s using standard technology that has been developed and tested by the Petersburg Nuclear Physics Institute. The characteristics of such detectors were described in [39, 40, 41, 42].
@@
The removable square-shaped transmission detector has a thickness of . The energy resolution of the cooled detector was measured to be FWHM = 7.9 keV. Switching the transmission detector into anti-coincidence mode with the Si(Li)- detector allows one to determine the - and X-ray contribution to the Si(Li) detector spectrum [40].
The 144Ce source in the form of a dried drop of colloidal solution was placed on the surface of a mylar substrate with thickness of m. The substrate was fixed in a delrin ring and placed on a supporting structure at the distance of mm from the detector surface. The thickness distribution over the spot of the dried source 144Ce was determined from the measured energy spectrum of alpha particles originating from the decays of 241Am and 244Cm, which were preserved in the source as low activity impurities after purification of the 144Ce source. The average thickness of the source was determined as m.
A HPGe-detector 160 volume in a low background setup was used to detect additional -activity in the -source. The main extraneous ’s are associated with the isotope activity, which was 0.004 relative to the activity at the end of the measurements. The comparison was made using -lines with energies of 723 keV and 697 keV for and , respectively. The next most important activity of is an order of magnitude less than the activity of .
The entire installation of the Si(Li)-detector and source was placed in a vacuum cryostat and was cooled down to the temperature of liquid nitrogen (LN). The Si(Li)-detector was equipped with a charge-sensitive preamplifier with a field-effect transistor placed inside the vacuum cryostat and cooled down to the LN temperature. The negative bias voltage of kV was applied directly to the gold coating of the detector.
@@
In order to select transitions to the state, we set the condition on the energy of -quantum registered by BGO detector to be above 1 MeV. This condition is fulfilled for the -quanta from the full absorption peak and part of the Compton scattering tail and allows for exclusion of transitions to state followed by a single 697 keV -quantum. The energy resolution of the BGO detector allows separating reliably the energies of 697 keV and 1 MeV.
The spectrum of Si(Li) detectors in coincidence with BGO signals that exceeded 1 MeV consists of the -spectrum of the allowed transition of to the excited state of and random coincidences. As will be shown below the spectrum shape is in good accordance with theoretical shape of the allowed -transition with 812 keV endpoint. The choice of relatively small BGO-detector was driven by requirement for high -registration efficiency in case of low background of random coincidences. The agreement of the total measured spectrum shape to the shape of the allowed -transition represents an important validity criterion for the measurements, response function and fitting procedure.
The energy calibration was performed with help of total absorption peaks for gamma quanta with energies of keV and keV of the 207Bi source (Fig. 2, [39]). The energy resolution FWHM (full width at half maximum), determined by the conversion electron peak with an energy of keV, was keV ( eV). The results of Monte Carlo simulation of the calibration spectrum show good agreement with the measured spectrum, including the electron backscattering, edges and shape of Compton scattering [40].
The spectrum of measured by Si(Li)-detector in coincidence with transition detector and the spectrum simulated with Monte Carlo method are shown in Fig. 3 [40]. When passing through a thin 300 m detector, conversion electrons lose energy in accordance with the Landau distribution and form two broad peaks in the Si(Li)-spectrum, shifted by an energy of about 100 keV. The lower plot shows the full absorption detector spectrum obtained without transition detector. The Si(Li)-detector response functions calculated by Monte Carlo method using the real ”target - detector” geometry of for electrons with different energies are also shown.
Measurements were carried out continuously for 2024 hours, with short technological stops. The long duration of the measurements was defined by the low activity of the 144Ce source, which was Bk. The drift of the gain coefficient of the Si(Li) detector spectrometric channel during the long-term measurements was compensated by re-calibrations driven by the peak of conversion electrons with an energy of 91.5 keV at decays of 144Ce and the 5.902 MeV -peak of 244Cm, which were present in the measured spectra.
@@
The original spectrometer with 4 geometry was designed on the basis of two Si(Li) detectors with the thicknesses of about mm each, which exceeds the ionisation path of electrons with an energy of MeV. The detectors were specially manufactured for this experiment and they were made in the “top-hat” shape with external diameters of mm and mm, height of mm and diameters of the sensitive area of mm and mm. The Si(Li) detectors had different external diameters for the convenience of assembling the spectrometer with two detectors docked closely (Fig. 1, (B)) [42, 43, 44, 45].
@@
The total registered energy spectrum is obtained by summing individual spectra from each of the Si(Li)-detectors and the energy spectrum recorded by both detectors in coincidence. The obtained spectrum solves the response function problem associated with backscattering of electrons from the detector surface.
@@ III DATA ANALYSIS
The peculiarity of the described above -spectrometer is that it registers with high efficiency the IB photons with energy up to 100 keV, which are considered as part of the radiative corrections. For example, 1 MeV electron appearing in -decay emits % of the IB photons with an energy of less than keV, of which only % leave the -detector without interaction. As a result, we used two different forms of the radiative correction function fitting the data of “target-detector” and -spectrometer. For the -spectrometer we used the correction introduced for the neutrino spectrum [55, 56] with the natural replacement of by .
@@ III.1 Analysis of the data of spectrometer “target-detector” type
The nuclear structure shape factor is the function searched in the problem of the -spectrum measurement. The shape factor was parameterized in powers of similarly to [58]:
where and are free parameters. As a result of the analysis, the following expression for the shape factor function was obtained:
The accuracy of the obtained result (3) is limited by the weak activity of the 144Ce source, which requires a small distance from the detector to the target, that leads to large angles of incidence of electrons onto the crystal surface and an increased proportion of backscattered electrons. The accuracy in determination of the the nuclear shape factor parameters and could be significantly increased by usage of a more active source with “target-detector” spectrometer.
@@ III.2 Analysis of the data of 4-spectrometer
Although the response of the spectrometer is close to the Gaussian function, one should take into account the passage of the insensitive layers of the source and detectors. In order to account for these effects detailed Monte Carlo calculations were carried out for different electron energies.
The detector response function determined as a result of fitting looks like as a convolution of a Gaussian function with an exponential tail, which describes small energy losses in the target and non-sensitive layers of the detectors. As an example, for 1 MeV electrons the standard deviation for response function (the square root of the second central moment ) is 56.8 keV and the fraction of electrons detected with an energy less than 0.850 MeV is 3.4%.
The response function calculated by the Monte Carlo method was described analytically as:
@@
As a result of the analysis, the following values of the parameters and of the nuclear shape factor function were obtained:
@@
The -spectrum has been previously measured in several works [58, 59, 60, 61, 62] that used different expressions for the Fermi function and for the shape factor. Therefore, a direct comparison of the results obtained is difficult. Nevertheless, our result for is more precise but consistent with [62, 61] within experimental uncertainties.
@@ IV NEUTRINO SPECTRUM
Fig. 8 illustrates the spectrum of electron antineutrinos from 144Pr source. The numerical values for neutrinos with energies above the IBD threshold are shown in the Tab. 1. The uncertainties of are calculated with toy Monte Carlo method based on the uncertainties of the evaluated nuclear shape factor parameters and from (5) and taking into account that these parameters are in quite strong anticorrelation with correlation coefficient of 0.973. The parameters evaluated with this method were simulated 10000 times and used for performing 2-dimensional spectrum density plot. The RMS of density distributions of its monoenergetic projections was taken as the statistical uncertainty of the spectrum.
@@
Neutrinos from the decay of have the same intensity as -neutrinos, but their energies are less than 318 keV (the region to the left of line 1 in Fig. 8). The threshold value of the IBD reaction on hydrogen in the laboratory system is (line 2).
@@
Considering the precision given by our measurement one could apply toy Monte Carlo to convert the uncertainties on the shape factor parameters to the fraction of the -spectrum above the threshold IBD reaction that is , so the statistic precision of this value is as low as 0.012 %. There are two main causes of systematic errors. The first uncertainty evaluated as the difference of the values derived for two boundary cases of low-energy screening correction is 0.00022 (0.044 %). A more significant systematic error is associated with the 1.2 keV uncertainty of the energy scale at 1.806 MeV which results in a value of 0.00065 (0.13 %).
The total cross section of the IBD reaction is determined by the integral from the threshold to the maximum neutrino energy over the product of the IBD reaction cross section and the neutrino flux (Fig.9):
@@
Using the cross section of IBD reaction from [63, 64, 65, 66], one could evaluate the integrated cross section, that is cm2 per one decay of nucleus. The total systematic error reported here is obtained by taking into account the uncertainty in the screening correction (=0.0041 (0.09%)) and inaccuracy of the energy scale (=0.012 (0.25%)). The latter can be substantially improved by any refinement of the -value of -decay. The obtained precision of in % would correspond to an experimental uncertainty of neutrino count rate expected for a sterile neutrino experiment using Pr source. In this case this value would be sub-dominated by other experimental uncertainties, e.g. the precision of the source calorimetry for Borexino-SOX experiment has reached precision of % [38].
@@
We can also provide an analytical interpolation of the spectrum, that could be expressed in two energy intervals. For neutrino energies above the unique forbidden transition endpoint of MeV, the spectrum is interpolated as follows ( in MeV units):
@@ V CONCLUSION
The -spectra of the source were studied using two spectrometers based on semiconductor Si(Li) detectors. The spectra were measured using a beta spectrometer in the classical “target-detector” scheme and with an original full absorption spectrometer consisting of two of Si(Li) detectors. The models of the beta spectrometer response function calculated through the Monte Carlo simulation were tested by fitting the spectrum of the allowed beta transition and have shown good agreement between the response models and experimental data. The function of the nuclear shape factor of the ground state beta transition in has been defined as . The electron antineutrino spectrum of the decays was obtained from the measured -spectrum and the shape factor parameters. The reduced cross section for the inverse beta decay reaction for the electron anti-neutrino source is cm2 per which is the most accurate measurement up to date.